# Optimal time estimation and the clock uncertainty relation for stochastic processes

### v1

@@ III.1 Counting observables and the BLUE
where ${\rm Cov}[A,B]={\rm E}[AB]-{\rm E}[A]{\rm E}[B]$. Here, $J_{\mu\sigma}={\rm E}[I_{\mu\sigma}]$ is the mean steady-state current (i.e. the mean number of transitions per unit time from $\sigma\to\mu$), while $D_{\alpha\beta\mu\sigma}$ is the diffusion tensor. Using the techniques of Ref. [44], we show in Appendix B that
@@
$D=\sum_{\alpha,\beta,\mu,\sigma=1}^{d}w_{\alpha\beta}D_{\alpha\beta\mu\sigma}w_{\mu\sigma}\equiv\vec{w}\cdot\mathbbm{D}\cdot\vec{w}.$ (26)
Here, $\vec{w}$ and $\vec{J}$ denote “vectors” with $d^{2}$ components, while the tensor $\mathbbm{D}$ maps one such vector to another.
@@
$\mathcal{S}=\frac{J^{2}}{D}=\frac{\left(\vec{w}\cdot\vec{J}\;\right)^{2}}{\vec{w}\cdot\mathbbm{D}\cdot\vec{w}}.$ (27)
A current that maximises the SNR is known as a hyperaccurate current [66, 67]; here, we extend this notion to general counting observables. Since $\mathcal{S}$ is invariant under uniform rescaling of the weights, $\vec{w}\to r\vec{w}$, we are free to rescale $\vec{w}$ to satisfy $\vec{w}\cdot\vec{J}=1$. This guarantees an unbiased estimator [cf. Eqs. (6) and (25)]. Maximising $\mathcal{S}$ is then equivalent to minimising the denominator of Eq. (27) under the constraint $\vec{w}\cdot\vec{J}=1$, which yields a condition that the optimal weights must satisfy (see Appendix B):
$\frac{\mathbbm{D}\cdot\vec{w}}{\vec{w}\cdot\mathbbm{D}\cdot\vec{w}}=\vec{J}.$ (28)
If $\mathbbm{D}$ were invertible, the solution of this equation would be of the form $\vec{w}\propto\mathbbm{D}^{-1}\cdot\vec{J}$, which gives more weight to elementary currents with low variance relative to their mean, while also accounting for current-current correlations. In general, however, the diffusion tensor is singular: it possesses at least $d-1$ linearly independent, null eigenvectors $\vec{w}$ such that $\mathbbm{D}\cdot\vec{w}=0$ (see Appendix B for their explicit construction). Therefore, the solution of Eq. (28) is not unique, and there exist several inequivalent counting observables with maximal SNR.
A particularly elegant solution of Eq. (28) is $w_{\mu\sigma}=\Gamma_{\sigma}^{-1}$, as can be verified directly (see Appendix B for details). The resulting time estimator is
@@
as shown in Appendix B. The variance then follows as
@@
and hence Ineq. (20) is satisfied for all $t$. Note that this implies that the variance of $\Theta(\mathbf{n})$ [Eq. (16)] may drop below that of the BLUE at intermediate times, as illustrated in Fig. 4 and discussed in more detail in Appendix F.
@@ III.2 Comparison to other precision bounds
TUR $\dfrac{\dot{\Sigma}}{2}=\displaystyle\frac{1}{2}\sum_{\mu,\sigma}p_{\sigma}^{\rm ss}R_{\mu\sigma}\ln\left(\frac{R_{\mu\sigma}}{R_{\sigma\mu}}\right)$ $\begin{array}[]{c}\Theta(\bm{\sigma})=\vec{w}\cdot\vec{N}\\ w_{\mu\sigma}=-w_{\sigma\mu}\end{array}$
@@
These findings are consistent with previous studies showing that precision is bounded tightly by dissipation close to equilibrium and by activity far away from it [19, 29]. Yet, in addition, Fig. 5(b) reveals that the relevant measure of activity far from equilibrium is the inverse mean residual time, $\mathcal{T}^{-1}$, as this generically yields the tightest bound on current fluctuations within this regime.
@@ III.3 Relation between the KUR and CUR
$\mathcal{F}_{\lambda}[P(\bm{\sigma},\mathbf{t}|t,\lambda)]=-{\rm E}\left[\partial_{\lambda}^{2}\ln P(\bm{\sigma},\mathbf{t}|t,\lambda)\big{|}t,\lambda\right]=\frac{{\rm E}[N|t,\lambda]}{1+\lambda}.$ (35)
@@
This can be shown by introducing new integration variables $t_{j}^{\prime}=(1+\lambda)t_{j}$ in Eq. (33), as detailed in Appendix H. For counting observables, or any other observable $\Theta(\bm{\sigma})$ that does not depend explicitly on the waiting times $\bm{\tau}$, the second term on the right-hand side of Eq. (36) vanishes. Therefore, combining Eqs. (34)–(36), setting $\lambda=0$, and rearranging, we obtain the KUR (valid for all $t>0$[19]
@@
In fact, we show in Appendix H that
@@
Figure 6: Asymptotic signal-to-noise ratio $\mathcal{S}$ of the observable $\Theta(\bm{\sigma},\bm{\tau})=\sum_{j=0}^{N}\Gamma_{\sigma_{j}}\tau_{j}$ computed for $10^{4}$ randomly generated, fully connected stochastic networks (see Appendix H for details of the calculation). We use the sampling procedure described in the caption of Fig. 5, with connectivity $c=1$ and dimension $d=10$.
@@
Finally, one may wonder whether it is possible to formulate the time estimation problem for this more general class of observables that depend explicitly on the waiting times. Interestingly, the CRB for time estimation does not exist in this case: the support of the distribution $P(\bm{\sigma},\mathbf{t}|t)$ depends on $t$ and so the corresponding Fisher information $\mathcal{F}_{t}[P(\bm{\sigma},\mathbf{t}|t)]$ cannot be defined [49]. Naturally, therefore, the problem of stochastic time estimation is only well posed if a clock is not already available to record the jump times $\mathbf{t}$, so that only the information in the marginal distribution $P(\bm{\sigma}|t)$ is accessible.
@@ IV.1 Resolution
Table 2: Three unbiased time estimators that are counting observables of the form $\nolinebreak{\Theta=\vec{w}\cdot\vec{N}}$. The properties of the Erlang estimator quoted above hold only for ring clocks, with arbitrary escape rates but only clockwise jumps allowed (see Appendix I).
@@
with $P(\sigma\to\mu|\aleph)$ the probability for the next monitored jump to be $\sigma\to\mu$. Eq. (43) describes the mean duration between recorded jumps, as defined by the estimator itself. A simple application of Bayes’ rule yields $P(\sigma\to\mu|\aleph)=J_{\mu\sigma}/\sum_{(\sigma\to\mu)\in\aleph}J_{\mu\sigma}$ (see Appendix C), and then using the unbiased constraint, $\vec{w}\cdot\vec{J}=1$, we obtain simply
@@
To demonstrate how the choice of estimator affects resolution, we consider the so-called Erlang clock [79, 80]. This is a ring network with $d$ states and fully biased transition rates, $R_{\mu\sigma}=\Gamma_{\sigma}\delta_{\mu,\sigma+1}$, such that only jumps in the clockwise direction are allowed (Fig. 7 inset). This model is exactly solvable in the homogeneous case, $\Gamma_{\sigma}=\Gamma$, as detailed in Appendix I. While the BLUE (29) weights every transition equally, $w_{\mu\sigma}=1/\Gamma$, one may also consider the Erlang estimator, where only the transition $d\to 1$ is counted, i.e. $w_{1d}=d/\Gamma$ and all other $w_{\mu\sigma}=0$. This is analogous to measuring time by counting seconds (single jumps) or minutes (complete revolutions around the ring). Both estimators are unbiased, ${\rm E}[\Theta]=t$, and have asymptotic variance scaling as ${\rm Var}[\Theta]/t\to\Gamma^{-1}=\mathcal{T}$, so they both saturate the CUR (see Appendix I).
@@
However, the two estimators yield significantly different tick statistics, as evidenced by their current autocorrelation functions, $F(\tau)={\rm Cov}[I(t+\tau),I(t)]$. A surprising property of the BLUE is that its current fluctuations are delta-correlated in time (see Eq. (30)). This property is characteristic of white or Poissonian noise, and it contradicts the intuitive expectation that the ticks of a reliable clock should be anti-bunched in time, which would imply that $F(\tau)>\lim_{\epsilon\to 0}F(\epsilon)$ for all $\tau>0$ [44]. By contrast, the autocorrelation function of the Erlang estimator does show this anti-bunched character, as shown in Fig. 7. This apparent regularity comes at the cost of decreased resolution: the resolution of the Erlang estimator is $\nu_{\rm Erl}=\Gamma/d$ while the BLUE has $\nu_{\rm BLUE}=\Gamma$ (see Appendix I). Indeed, the BLUE ticks after every jump, whereas each tick of the Erlang estimator requires the clock to complete $d$ jumps around the ring. Nevertheless, sacrificing resolution does not improve precision as measured by the SNR. In fact, at short times the Erlang estimator has a significantly larger variance than the BLUE. In Appendix I we show that $\lim_{t\to 0}{\rm Var}[\Theta_{\rm Erl}]/t=d/\Gamma$, while for the BLUE we have ${\rm Var}[\Theta_{\rm BLUE}]/t=\Gamma^{-1}$ at all times.
To illustrate how the variance of the Erlang estimator scales at intermediate times, we return to the two-state system with unbalanced rates in Fig. 4. In the case of inhomogeneous rates $\Gamma_{\sigma}$, the Erlang estimator is defined by weighting the transition $d\to 1$ as $w_{1d}=J_{1d}^{-1}$, setting all other $w_{\mu\sigma}=0$. This yields an unbiased estimator with minimal variance at long times, $\lim_{t\to\infty}{\rm Var}[\Theta_{\rm Erl}]/t=\mathcal{T}$, and resolution $\nu_{\rm Erl}=\mathcal{A}/d$ that is $d$ times smaller than the BLUE. These expressions hold for any $d$, so long as only clockwise transitions are allowed; see Table 2 for a summary and Appendix I for details. Figure 4 shows the scaled variance of this estimator for $d=2$ as a function of time (black squares), demonstrating that the expected error is persistently larger than the variance of the BLUE (upper dashed line) while tending to the same value as $t\to\infty$. Overall, we conclude that maximising both resolution and precision requires the estimator to count all observable events.

### v2

@@ III.1 Counting observables and the BLUE
where ${\rm Cov}[A,B]={\rm E}[AB]-{\rm E}[A]{\rm E}[B]$. Here, $J_{\mu\sigma}={\rm E}[I_{\mu\sigma}]$ is the mean steady-state current (i.e. the mean number of transitions per unit time from $\sigma\to\mu$), while $D_{\alpha\beta\mu\sigma}$ is the diffusion tensor. Using the techniques of Ref. [46], we show in Appendix G that
@@
$D=\sum_{\alpha,\beta,\mu,\sigma=1}^{d}w_{\alpha\beta}D_{\alpha\beta\mu\sigma}w_{\mu\sigma}\equiv\vec{w}\cdot\mathbb{D}\cdot\vec{w}.$ (26)
Here, $\vec{w}$ and $\vec{J}$ denote “vectors” with $d^{2}$ components, while the tensor $\mathbb{D}$ maps one such vector to another.
@@
$\mathcal{S}=\frac{J^{2}}{D}=\frac{\left(\vec{w}\cdot\vec{J}\;\right)^{2}}{\vec{w}\cdot\mathbb{D}\cdot\vec{w}}.$ (27)
A current that maximises the SNR is known as a hyperaccurate current [75, 76]; here, we extend this notion to general counting observables. Since $\mathcal{S}$ is invariant under uniform rescaling of the weights, $\vec{w}\to r\vec{w}$, we are free to rescale $\vec{w}$ to satisfy $\vec{w}\cdot\vec{J}=1$. This guarantees an unbiased estimator [cf. Eqs. (6) and (25)]. Maximising $\mathcal{S}$ is then equivalent to minimising the denominator of Eq. (27) under the constraint $\vec{w}\cdot\vec{J}=1$, which yields a condition that the optimal weights must satisfy (see Appendix G):
$\frac{\mathbb{D}\cdot\vec{w}}{\vec{w}\cdot\mathbb{D}\cdot\vec{w}}=\vec{J}.$ (28)
If $\mathbb{D}$ were invertible, the solution of this equation would be of the form $\vec{w}\propto\mathbb{D}^{-1}\cdot\vec{J}$, which gives more weight to elementary currents with low variance relative to their mean, while also accounting for current-current correlations. In general, however, the diffusion tensor is singular: it possesses at least $d-1$ linearly independent, null eigenvectors $\vec{w}$ such that $\mathbb{D}\cdot\vec{w}=0$ (see Appendix G for their explicit construction). Therefore, the solution of Eq. (28) is not unique, and there exist several inequivalent counting observables with maximal SNR.
A particularly elegant solution of Eq. (28) is $w_{\mu\sigma}=\Gamma_{\sigma}^{-1}$, as can be verified directly (see Appendix G for details). The resulting time estimator is
@@
as shown in Appendix G. The variance then follows as
@@
and hence Ineq. (20) is satisfied for all $t$. Note that this implies that the variance of $\Theta(\mathbf{n})$ [Eq. (16)] may drop below that of the BLUE at intermediate times, as illustrated in Fig. 4 and discussed in more detail in Appendix E.
@@ III.2 Comparison to other precision bounds
TUR $\dfrac{\dot{\Sigma}}{2}=\displaystyle\frac{1}{2}\sum_{\mu,\sigma}J_{\mu\sigma}\ln\left(\frac{J_{\mu\sigma}}{J_{\sigma\mu}}\right)$ $\begin{array}[]{c}\Theta(\bm{\sigma})=\vec{w}\cdot\vec{N}\\ w_{\mu\sigma}=-w_{\sigma\mu}\end{array}$
@@
These findings are consistent with previous studies showing that precision is bounded tightly by dissipation close to equilibrium and by activity far away from it [19, 29]. Yet, in addition, Fig. 5(b) reveals that the relevant measure of activity far from equilibrium is the inverse mean residual time, $\mathcal{T}^{-1}$, as this generically yields the tightest bound on current fluctuations within this regime. In Appendix H, we also compare the CUR with two entropic bounds on counting observables derived recently in Ref. [40], finding again that the CUR is the tightest bound far from equilibrium.
@@ III.3 Relation between the KUR and CUR
$\mathcal{F}_{\lambda}[P(\bm{\sigma},\mathbf{t}|t,\lambda)]=-{\rm E}\left[\partial_{\lambda}^{2}\ln P(\bm{\sigma},\mathbf{t}|t,\lambda)\big{|}t,\lambda\right]=\frac{{\rm E}[N|t,\lambda]}{(1+\lambda)^{2}}.$ (35)
@@
This can be shown by introducing new integration variables $t_{j}^{\prime}=(1+\lambda)t_{j}$ in Eq. (33), as detailed in Appendix I. For counting observables, or any other observable $\Theta(\bm{\sigma})$ that does not depend explicitly on the waiting times $\bm{\tau}$, the second term on the right-hand side of Eq. (36) vanishes. Therefore, combining Eqs. (34)–(36), setting $\lambda=0$, and rearranging, we obtain the KUR (valid for all $t>0$[19]
@@
In fact, we show in Appendix I that
@@
Figure 6: Asymptotic signal-to-noise ratio $\mathcal{S}$ of the observable $\Theta(\bm{\sigma},\bm{\tau})=\sum_{j=0}^{N}\Gamma_{\sigma_{j}}\tau_{j}$ computed for $10^{4}$ randomly generated, fully connected stochastic networks (see Appendix I for details of the calculation). We use the sampling procedure described in the caption of Fig. 5, with connectivity $c=1$ and dimension $d=10$.
@@
Finally, one may wonder whether it is possible to formulate the time estimation problem for this more general class of observables that depend explicitly on the waiting times. Interestingly, the CRB for time estimation does not exist in this case: the support of the distribution $P(\bm{\sigma},\mathbf{t}|t)$ depends on $t$ and so the regularity condition ${\rm E}[\partial_{t}\ln P(\bm{\sigma},\mathbf{t}|t)]=0$, which is necessary to derive the CRB [51], does not hold. Naturally, therefore, the problem of stochastic time estimation is only well posed if a clock is not already available to record the jump times $\mathbf{t}$, so that only the information in the marginal distribution $P(\bm{\sigma}|t)$ is accessible.
III.4 Testing the CUR to detect hidden transitions
Figure 7: Detecting hidden transitions. We consider the hidden Markov model shown in the inset, with (hidden) transition rates $R_{\mu\sigma}=r_{\sigma}\delta_{\mu,\sigma+1}+{\rm e}^{-\Delta}r_{\mu}\delta_{\mu,\sigma-1}$ and periodic boundary conditions. Each data point corresponds to one random realisation generated by independently sampling the rates $r_{\sigma}\in[0,1]$ and bias $\Delta\in[0,5]$ from uniform distributions. We compute the optimal SNR $\tilde{\mathcal{S}}_{\rm max}$ and compare to the mean residual time $\tilde{\mathcal{T}}$ and waiting time $\tilde{\mathcal{A}}^{-1}$ for observable transitions $\mathtt{X}\leftrightarrow\mathtt{Y}$; see Appendix J for details of the calculation. Hidden transitions are detected whenever $\tilde{\mathcal{S}}_{\rm max}\tilde{\mathcal{T}}\neq 1$. Hidden transitions are also witnessed if $\tilde{\mathcal{T}}<\tilde{\mathcal{A}}^{-1}$ (blue circles), or by violations of the KUR (top-left yellow quadrant), CUR (bottom-right blue quadrant), or both (top-right grey quadrant).
We have shown in Sec. III.1 that the CUR (1) is saturated by a class of optimal counting observables with maximal SNR. For such observables, the CUR becomes an equality
$\mathcal{S}_{\rm max}=\mathcal{T}^{-1}.$ (43)
If all transitions are observable, this relation can be verified experimentally from the measured jump statistics, as explained below. Conversely, a violation of Eq. (43) would indicate that the observed dynamics fails to satisfy one of the conditions for the validity of the CUR. To recap, these conditions are that the system is statistically stationary and undergoes classical, autonomous, Markovian dynamics as described by a time-homogeneous master equation $\partial_{t}\mathbf{p}=\mathsf{L}\mathbf{p}$. While these assumptions hold microscopically in a broad range of biological, chemical, and mesoscopic scenarios, often only a subset of the system’s states or transitions can be observed directly. The observable degrees of freedom are coupled to the remaining, hidden states, resulting in effectively non-Markovian dynamics (a so-called hidden Markov model). The presence and influence of such hidden states can thus be detected through violations of the CUR. Crucially, this approach works in the steady state, thereby complementing the recent proposal of Ref. [53] that is restricted to non-stationary dynamics.
To illustrate the general idea, consider a system switching between two observable “meso-states” $\mathtt{X}$ and $\mathtt{Y}$, each of which may comprise numerous microscopic states [85]; an example is illustrated in the inset of Fig. 7. This scenario could describe a cellular chemoreceptor with “bound” and “unbound” states [86, 78], a gene flipping between “on” and “off” transcription states [87], or a molecular motor switching between distinct motional states [88, 89]. Switching between the meso-states manifests as a telegraph process, which appears non-Markovian since the waiting times between jumps depend on the inaccessible micro-state before and after each jump.
To test the relation (43), an experimenter must use their observations to construct the maximum SNR and the mean residual time. The latter can be inferred in various ways from the empirical waiting-time statistics (in this section only, we assume the experimentalist has a reliable external clock). For example, if $\tau$ denotes the waiting time between any pair of consecutive transitions, then [65, 90] (see also Appendix C)
$\tilde{\mathcal{T}}=\frac{{\rm E}[\tau^{2}]}{2{\rm E}[\tau]},$ (44)
which holds for an arbitrary (non-Markovian) stationary process. Here, we use a tilde to distinguish an observable quantity $\tilde{\mathcal{T}}$ from its “true” value $\mathcal{T}$ pertaining to the hidden Markov network. Clearly, $\tilde{\mathcal{T}}\geq\mathcal{T}$ because some transitions are missed, and so the inequality $\mathcal{S}\leq\tilde{\mathcal{T}}^{-1}$ can be violated if the SNR is large enough.
To increase the likelihood of detecting such a violation, our experimenter should find the maximum SNR attainable from the empirical jump statistics, $\tilde{\mathcal{S}}_{\rm max}$. Let $N_{\mathtt{XY}}$ denote the number of transitions $\mathtt{Y}\to\mathtt{X}$ (shown as pink arrows in the inset of Fig. 7) and $N_{\mathtt{Y}\mathtt{X}}$ is the corresponding number of jumps $\mathtt{X}\to\mathtt{Y}$ (light blue arrows) observed over a duration $t$. Assuming stationary statistics and a large observation time, one constructs the mean currents $\tilde{\vec{J}}$ and diffusion tensor $\tilde{\mathbb{D}}$ as [c.f. Eq. (III.1)]
$\tilde{\vec{J}}=\lim_{t\to\infty}\frac{1}{t}\begin{pmatrix}{\rm E}[N_{\mathtt{XY}}]\\ {\rm E}[N_{\mathtt{Y}\mathtt{X}}]\end{pmatrix},$ (45)
$\tilde{\mathbb{D}}=\lim_{t\to\infty}\frac{1}{t}\begin{pmatrix}{\rm Var}[N_{\mathtt{XY}}]&{\rm Cov}[N_{\mathtt{XY}},N_{\mathtt{YX}}]\\ {\rm Cov}[N_{\mathtt{YX}},N_{\mathtt{XY}}]&{\rm Var}[N_{\mathtt{YX}}]\end{pmatrix}.$ (46)
Following the same logic as Sec. III.1 and Appendix G, the maximal SNR is then given by [c.f. Eq. (154)]
$\tilde{\mathcal{S}}_{\rm max}=\tilde{\vec{J}}\cdot\tilde{\mathbb{D}}^{+}\cdot\tilde{\vec{J}}.$ (47)
This procedure directly generalises to an arbitrary number of observable meso-states. However, in the present case of two meso-states, we show in Appendix J that $N_{\mathtt{X}\mathtt{Y}}$ and $N_{\mathtt{Y}\mathtt{X}}$ are perfectly correlated at long times. Without loss of generality, therefore, it suffices to measure only one of them, say $N_{\mathtt{X}\mathtt{Y}}$, to obtain $\tilde{\mathcal{S}}_{\rm max}=\lim_{t\to\infty}{\rm E}[N_{\mathtt{XY}}]^{2}/(t{\rm Var}[N_{\mathtt{XY}}])$.
Figure 7 displays a numerical example inspired by chemoreception models of Refs. [86, 58, 78]. We randomly sample transition rates between the microscopic states of the hidden Markov model, including a random bias favouring jumps in the clockwise direction to break detailed balance. The maximal SNR is reduced by the omission of unobserved jumps, while the mean residual time is increased. However, they are generally changed by different amounts, so that $\tilde{\mathcal{S}}_{\rm max}$ may be either larger or smaller than $\tilde{\mathcal{\mathcal{T}}}^{-1}$. Whenever $\tilde{\mathcal{S}}_{\rm max}\tilde{\mathcal{\mathcal{T}}}\neq 1$, therefore, the influence of hidden transitions is detected. We observe in Fig. 7 that this is generically the case, except for in scenarios of high symmetry that arise when the bias is small so that detailed balance is approximately restored.
In principle, non-Markovian dynamics could also be witnessed by a violation of the KUR bound $\tilde{\mathcal{S}}_{\rm max}\leq\tilde{\mathcal{A}}$, where $\tilde{\mathcal{A}}$ is the mean rate of observed transitions. This is an imperfect witness, however, because the KUR bound is generally not achievable even for perfectly Markovian networks, as shown in Sec. III.3. To illustrate this, we also show the ratio $\tilde{\mathcal{S}}_{\rm max}/\tilde{\mathcal{A}}$ in Fig. 7. Only for those points in the upper-left (yellow) quadrant of Fig. 7 is the KUR apparently violated while the CUR is not. In such cases, however, one could also detect non-Markovianity by simply noting that $\tilde{\mathcal{A}}\tilde{\mathcal{T}}<1$. This can be tested directly from the mean waiting time, $\tilde{\mathcal{A}}^{-1}$, and its variance: from Eq. (44) we thus have
$\tilde{\mathcal{A}}\tilde{\mathcal{T}}=\frac{1}{2}\left(1+\frac{{\rm Var}[\tau]}{{\rm E}[\tau]^{2}}\right).$ (48)
A Markov jump process has ${\rm Var}[\tau]\geq{\rm E}[\tau]^{2}$ (see Appendix C), while this condition is generally violated for a non-Markovian process, as indicated by the blue points in Fig. 7. By comparison, the CUR saturation condition (43) provides the most stringent test of Markovian dynamics, as it is violated by all points away from the centre of Fig. 7. Observation of such violations could also be used to detect the presence of quantum-coherent or non-autonomous dynamics, as these are also inconsistent with the CUR.
One may envision other tests of classical Markovianity based on the waiting times, e.g. by confirming that they are (conditionally) exponentially distributed. This would require reconstruction and hypothesis testing of conditional waiting-time distributions such as $W(\tau|\mathtt{X})$ and $W(\tau|\mathtt{Y})$, the distribution of waiting time $\tau$ after a jump is observed into state $\mathtt{X}$ or $\mathtt{Y}$, respectively. On the contrary, testing the CUR requires only the mean and variance of the unconditional waiting times to evaluate Eq. (44). Our approach is thus comparable, yet provides complementary information, to recent proposals for inferring entropy production from conditional waiting-time statistics [58, 59, 60]. In general, entropy is produced both by observable transitions and by unobserved dynamics within the meso-states [85]. Identifying the presence of hidden states can therefore help to elucidate these different contributions to the dissipation inferred from partial observations of a non-equilibrium system.
@@ IV.1 Resolution
Table 2: Three unbiased time estimators that are counting observables of the form $\nolinebreak{\Theta=\vec{w}\cdot\vec{N}}$. The properties of the Erlang estimator quoted above hold only for ring clocks, with arbitrary escape rates but only clockwise jumps allowed (see Appendix K).
@@
with $P(\sigma\to\mu|\aleph)$ the probability for a given monitored jump to be $\sigma\to\mu$. Eq. (49) describes the mean duration between recorded jumps, as defined by the estimator itself. Following the same logic used to obtain the jump steady state in Sec. II.2 and Appendix B, it is straightforward to show that $P(\sigma\to\mu|\aleph)=J_{\mu\sigma}/\sum_{(\sigma\to\mu)\in\aleph}J_{\mu\sigma}$. Then, using the unbiased constraint, $\vec{w}\cdot\vec{J}=1$, we obtain
@@
To demonstrate how the choice of estimator affects resolution, we consider the so-called Erlang clock [94, 6, 95]. This is a ring network with $d$ states and fully biased transition rates, $R_{\mu\sigma}=\Gamma_{\sigma}\delta_{\mu,\sigma+1}$, such that only jumps in the clockwise direction are allowed (Fig. 8 inset). This model is exactly solvable in the homogeneous case, $\Gamma_{\sigma}=\Gamma$, as detailed in Appendix K. While the BLUE (29) weights every transition equally, $w_{\mu\sigma}=1/\Gamma$, one may also consider the Erlang estimator, where only the transition $d\to 1$ is counted, i.e. $w_{1d}=d/\Gamma$ and all other $w_{\mu\sigma}=0$. This is analogous to measuring time by counting seconds (single jumps) or minutes (complete revolutions around the ring). Both estimators are unbiased, ${\rm E}[\Theta]=t$, and have asymptotic variance scaling as ${\rm Var}[\Theta]/t\to\Gamma^{-1}=\mathcal{T}$, so they both saturate the CUR (see Appendix K).
@@
However, the two estimators yield significantly different tick statistics, as evidenced by their current autocorrelation functions, $F(\tau)={\rm Cov}[I(t+\tau),I(t)]$. A surprising property of the BLUE is that its current fluctuations are delta-correlated in time (see Eq. (30)). This property is characteristic of white or Poissonian noise, and it contradicts the intuitive expectation that the ticks of a reliable clock should be anti-bunched in time, which would imply that $F(\tau)>\lim_{\epsilon\to 0}F(\epsilon)$ for all $\tau>0$ [46]. By contrast, the autocorrelation function of the Erlang estimator does show this anti-bunched character, as shown in Fig. 8. This apparent regularity comes at the cost of decreased resolution: the resolution of the Erlang estimator is $\nu_{\rm Erl}=\Gamma/d$ while the BLUE has $\nu_{\rm BLUE}=\Gamma$ (see Appendix K). Indeed, the BLUE ticks after every jump, whereas each tick of the Erlang estimator requires the clock to complete $d$ jumps around the ring. Nevertheless, sacrificing resolution does not improve precision as measured by the SNR. In fact, at short times the Erlang estimator has a significantly larger variance than the BLUE. In Appendix K we show that $\lim_{t\to 0}{\rm Var}[\Theta_{\rm Erl}]/t=d/\Gamma$, while for the BLUE we have ${\rm Var}[\Theta_{\rm BLUE}]/t=\Gamma^{-1}$ at all times.
To illustrate how the variance of the Erlang estimator scales at intermediate times, we return to the two-state system with unbalanced rates in Fig. 4. In the case of inhomogeneous rates $\Gamma_{\sigma}$, the Erlang estimator is defined by weighting the transition $d\to 1$ as $w_{1d}=J_{1d}^{-1}$, setting all other $w_{\mu\sigma}=0$. This yields an unbiased estimator with minimal variance at long times, $\lim_{t\to\infty}{\rm Var}[\Theta_{\rm Erl}]/t=\mathcal{T}$, and resolution $\nu_{\rm Erl}=\mathcal{A}/d$ that is $d$ times smaller than the BLUE. These expressions hold for any $d$, so long as only clockwise transitions are allowed; see Table 2 for a summary and Appendix K for details. Figure 4 shows the scaled variance of this estimator for $d=2$ as a function of time (black squares), demonstrating that the expected error is persistently larger than the variance of the BLUE (upper dashed line) while tending to the same value as $t\to\infty$. Overall, we conclude that maximising both resolution and precision requires the estimator to count all observable events.